At what effective rate of interest will the present value of $1,400 at the end of three years and $800 at the end of six years be equal to $1,500?
Answer #1

FV of both will equal

So 1400*(1+r)3 = 800*(1+r)6
14/8 = (1+r)3
r = (14/8)1/3-1 = 20.51%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well